# Dirac Sources for Nonmetricity and Torsion in Metric-affine Gravity

### v1

@@ 5.3 Field equations
$\mathscr{Q}_{\quad ab}^{0}$ $=$ $\frac{4\alpha}{\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\mu}\nu}&I_{\bar{\mu}\rho}&I_{\bar{\mu}\sigma}\\ I_{\bar{\mu}\nu}&0&0&0\\ I_{\bar{\mu}\rho}&0&0&0\\ I_{\bar{\mu}\sigma}&0&0&0\end{array}\right)$
$\mathscr{Q}_{\quad ab}^{1}$ $=$ $\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}2I_{\bar{\mu}\sigma}&I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0\\ I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}&-2I_{\bar{\nu}\rho}&0&I_{\bar{\mu}\nu}+I_{\bar{\rho}\sigma}\\ I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0&2I_{\bar{\nu}\rho}&I_{\bar{\mu}\rho}+I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\mu}\nu}+I_{\bar{\rho}\sigma}&I_{\bar{\mu}\rho}+I_{\bar{\nu}\sigma}&2I_{\bar{\mu}\sigma}\end{array}\right)$
$\mathscr{Q}_{\quad ab}^{2}$ $=$ $\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}-2R_{\bar{\mu}\sigma}&R_{\bar{\nu}\sigma}+R_{\bar{\mu}\rho}&R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}&2\left(D_{\bar{\mu}\mu}+D_{\bar{\sigma}\sigma}\right)\\ R_{\bar{\nu}\sigma}+R_{\bar{\mu}\rho}&-2R_{\bar{\nu}\rho}&2\left(D_{\bar{\nu}\nu}-D_{\bar{\rho}\rho}\right)&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}\\ R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}&2\left(D_{\bar{\nu}\nu}-D_{\bar{\rho}\rho}\right)&2R_{\bar{\nu}\rho}&R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}\\ 2\left(D_{\bar{\mu}\mu}+D_{\bar{\sigma}\sigma}\right)&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}&-2R_{\bar{\mu}\sigma}\end{array}\right)$
$\mathscr{Q}_{\quad ab}^{3}$ $=$ $-\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}-2I_{\bar{\mu}\rho}&I_{\bar{\nu}\rho}+I_{\bar{\mu}\sigma}&0&I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}\\ I_{\bar{\nu}\rho}+I_{\bar{\mu}\sigma}&-2I_{\bar{\nu}\sigma}&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&-2I_{\bar{\mu}\rho}&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}\\ I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}&2I_{\bar{\nu}\sigma}\end{array}\right)$
@@
6 Special cases
We gain some insight by looking at restricted spinors. For a spin-up electron at rest, $\psi$ reduces to $\psi_{e0}=\left(\mu,0,0,0\right)$. Then all components of the torsion vanish $\mathcal{T}_{\;\;\;ab}^{0}=\mathcal{T}_{\;\;\;ab}^{1}=\mathcal{T}_{\;\;\;ab}^{3}=0$ except for

### v2

@@ 5.3 Field equations
$\mathscr{Q}_{\quad ab}^{0}$ $=$ $\frac{4\alpha}{\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\mu}\nu}&I_{\bar{\mu}\rho}&I_{\bar{\mu}\sigma}\\ I_{\bar{\mu}\nu}&0&0&0\\ I_{\bar{\mu}\rho}&0&0&0\\ I_{\bar{\mu}\sigma}&0&0&0\end{array}\right)$ (62)
$\mathscr{Q}_{\quad ab}^{1}$ $=$ $\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}2I_{\bar{\mu}\sigma}&I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0\\ I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}&-2I_{\bar{\nu}\rho}&0&I_{\bar{\mu}\nu}+I_{\bar{\rho}\sigma}\\ I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0&2I_{\bar{\nu}\rho}&I_{\bar{\mu}\rho}+I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\mu}\nu}+I_{\bar{\rho}\sigma}&I_{\bar{\mu}\rho}+I_{\bar{\nu}\sigma}&2I_{\bar{\mu}\sigma}\end{array}\right)$ (67)
$\mathscr{Q}_{\quad ab}^{2}$ $=$ $\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}-2R_{\bar{\mu}\sigma}&R_{\bar{\nu}\sigma}+R_{\bar{\mu}\rho}&R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}&2\left(D_{\bar{\mu}\mu}+D_{\bar{\sigma}\sigma}\right)\\ R_{\bar{\nu}\sigma}+R_{\bar{\mu}\rho}&-2R_{\bar{\nu}\rho}&2\left(D_{\bar{\nu}\nu}-D_{\bar{\rho}\rho}\right)&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}\\ R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}&2\left(D_{\bar{\nu}\nu}-D_{\bar{\rho}\rho}\right)&2R_{\bar{\nu}\rho}&R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}\\ 2\left(D_{\bar{\mu}\mu}+D_{\bar{\sigma}\sigma}\right)&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}&-2R_{\bar{\mu}\sigma}\end{array}\right)$ (72)
$\mathscr{Q}_{\quad ab}^{3}$ $=$ $-\frac{2\alpha}{\kappa}\left(\begin{array}[]{cccc}-2I_{\bar{\mu}\rho}&I_{\bar{\nu}\rho}+I_{\bar{\mu}\sigma}&0&I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}\\ I_{\bar{\nu}\rho}+I_{\bar{\mu}\sigma}&-2I_{\bar{\nu}\sigma}&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&-2I_{\bar{\mu}\rho}&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}\\ I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}&2I_{\bar{\nu}\sigma}\end{array}\right)$ (77)
@@
6 Summary and special cases
We studied the coupling of metric-affine gravity as a $GL\left(4\right)$ gauge theory, with a Dirac spinor field. The principal difficulty to overcome is that $GL\left(4\right)$ has no natural spinor representation and does not preserve the Lorentz metric required to define one. The connection, however, is Lie algebra valued, and the Lie algebra $\mathfrak{gl}\left(4\right)$ is isomorphic to Clifford algebra $Cl(3,1)$. Both are isomorphic to the real form of the Clifford algebra $Cl\left(2,2\right)$, so we expanded the $\mathfrak{gl}\left(4\right)$-valued connection as a real linear combination of the $Cl\left(2,2\right)$ basis. Replacing $\tilde{\gamma}^{2}\rightarrow i\gamma^{2}$ where $\tilde{\gamma}^{2}$ lies in the $Cl\left(2,2\right)$ basis and $\gamma^{2}$ lies in the $Cl\left(3,1\right)$ basis then gave the expansion of the $\mathfrak{gl}\left(4\right)$ connection in terms of Dirac matrices. This allows coupling to the Dirac field.
To separate contributions to torsion from sources for nonmetricity we identified the $\mathfrak{so}\left(3,1\right)$ subalgebra within $\mathfrak{gl}\left(4\right)$. That subset of generators was identified with couplings to torsion, with the remainder of the connection coupling to nonmetricity.
To carry out the variation, we wrote the $GL\left(4\right)$ curvature, separating torsion and nonmetric dependence from the usual Riemannian curvature. Variation showed the usual vanishing of torsion in vacuum, while the nonmetricity reduced to those parts expressible as a Weyl geometry with torsion. Similarly, we separated the spinor action into vacuum, torsion, and nonmetricity contributions.
Variation leads to Eq.(31),
$-\kappa\delta b_{c\Delta_{a}}tr\left(\Gamma^{\Delta_{a}}\mathscr{T}^{c}\right)-\frac{\kappa}{4}\delta b_{c\Delta_{s}}tr\left(\Gamma^{\Delta_{s}}\mathscr{Q}_{c}\right)$ $=$ $\alpha Re\left(i\psi^{\dagger}h\gamma^{c}\delta b_{\Delta c}\hat{\Gamma}^{\Delta}\psi\right)$
where $tr\left(\Gamma^{\Delta_{a}}\mathscr{T}^{c}\right)$ and $tr\left(\Gamma^{\Delta_{s}}\mathscr{Q}_{c}\right)$ are traces over the gamma matrix expansions. These give linear combinations of certain torsion and nonmetricity components. The right side gives the corresponding combinations of Dirac currents.
Our main results, the Dirac sources for torsion and nonmetricity in metric-affine gravity, are presented as explicit matrices built from spinor components in Eqs.(57) and (77). This specificity is necessary because the usual covariant notation breaks down in many expressions. The 64 real scalars constructed from Dirac spinors are collected in the Appendix.
To conclude, we gain some insight by looking at restricted spinors. For a spin-up electron at rest, $\psi$ reduces to $\psi_{e0}=\left(\mu,0,0,0\right)$. Then all components of the torsion vanish $\mathcal{T}_{\;\;\;ab}^{0}=\mathcal{T}_{\;\;\;ab}^{1}=\mathcal{T}_{\;\;\;ab}^{3}=0$ except for
@@ Acknowledgements:
?refname?
  • [1] H. Weyl, “Gravitation and electricity,” Sitzungsber. Preuss. Akad. Wiss. Berlin (Math. Phys.) 1918 (1918) 465.
  • [2] Weyl, Hermann, Sitz. Königlich Preußischen Akademie Wiss. (1918) 465; H. Weyl, Ann. d. Physik (4) 59, (1919) 101; H. Weyl, Gött. Nachr. (1921) 99; H. Weyl, Raum, Zeit, Materie, Springer, Berlin, (1919-1923) .
  • [3] Hermann Weyl, Math. Zeitschr., 2 (1918b) 384.
  • [4] H. Weyl, “A new extension of relativity theory,” Annalen Phys. 364 (1919) 101 [Surveys High Energ. Phys. 5 (1986) 237] [Annalen Phys. 364 (1919) 101].
  • [5] É. Cartan, Sur une généralisation de la notion de courbure de Riemann et les espaces à torsion, C.R. Acad. Sci. (Paris) 174 (1922) 593–595; English translation by G.D. Kerlick: On a generalization of the notion of Riemann curvature and spaces with torsion, in: “Proc. of the 6th Course of Internat. School on Cosmology and Gravitation: Spin, Torsion, Rotation, and Supergravity” (Erice, 1979), Eds. P.G.Bergmann and V. De Sabbata (Plenum: New York, 1980) 489–491; with subsequent comments of A. Trautman, Comments on the paper by Élie Cartan: Sur une généralisation de la notion de courbure de Riemann et les espaces à torsion, pp. 493–496.
  • [6] E. Cartan, CR Acad. Sci. 174, p. 437, 1922.
  • [7] Cartan, Élie, Sur une généralisation de la notion de courbure de Riemann et les espaces à torsion, Comptes rendus de l’Académie des Sciences de Paris (in French) 174: (1922) 593–595.
  • [8] Cartan, Elie, Sur les variétés à connexion affine et la théorie de la relativité généralisée (première partie), Annales Scientifiques de l’École Normale Supérieure (in French), 40: (1923) 325–412. doi:10.24033/asens.751. ISSN 0012-9593.
  • [9] Cartan, Elie, Sur les variétés à connexion affine, et la théorie de la relativité généralisée (première partie) (Suite), Annales Scientifiques de l’École Normale Supérieure (in French). 41: (1924). 1–25. doi:10.24033/asens.753. ISSN 0012-9593.
  • [10] Einstein, Albert (1928). Riemann-Geometrie mit Aufrechterhaltung des Begriffes des Fernparallelismus, Preussische Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, 1928: 217–221.
  • [11] Kibble, T. W. B. (1961). "Lorentz Invariance and the Gravitational Field". Journal of Mathematical Physics. 2 (2): 212–221. Bibcode:1961JMP…..2..212K. doi:10.1063/1.1703702. ISSN 0022-2488. S2CID 54806287.
  • [12] Sciama, D. W., On the analogy between charge and spin in general relativity, Recent Developments in General Relativity. Warsaw: Polish Scientific Publishers, 1962., p.415
  • [13] Sciama, D. W. The Physical Structure of General Relativity, Reviews of Modern Physics. 36 (1): 463–469. (Jan 1, 1964). Bibcode:1964RvMP…36..463S. doi:10.1103/revmodphys.36.463. ISSN 0034-6861.
  • [14] F. W. Hehl and B. K. Datta, Nonlinear Spinor Equation and Asymmetric Connection in General Relativity, J. Math. Phys. 12, 1334 (1971); https://doi.org/10.1063/1.1665738
  • [15] K. Datta, Spinor fields in general relativity, II: Generalized field equations and application to the Dirac field, Il Nuovo Cimento B (1971-1996) volume 6, pages 16–28 (1971)
  • [16] F. W. Hehl, Spin and torsion in general relativity: I. Foundations, General Relativity and Gravitation volume 4, pages 333–349 (1973)
  • [17] Friedrich W. Hehl, Paul von der Heyde, and G. David Kerlick, General relativity with spin and torsion and its deviations from Einstein’s theory, Phys. Rev. D 10, (15 August 1974) 1066.
  • [18] Hehl, Friedrich W.; von der Heyde, Paul; Kerlick, G. David; Nester, James M. (1976-07-01). General relativity with spin and torsion: Foundations and prospects. Reviews of Modern Physics. 48 (3): 393–416. doi:10.1103/revmodphys.48.393
  • [19] Wheeler, James T., Weyl Geometry, General Relativity and Gravitation (2018) 50:80, arXiv:1801.03178 [gr-qc], Gen Relativ Gravit (2018) 50:80, DOI, https://doi.org/10.1007/s10714-018-2401-5.
  • [20] F.W. Hehl, G.D. Kerlick, P. Von Der Heyde, On a New Metric Affine Theory of Gravitation, Phys. Lett. B 63 (1976) 446-448
  • [21] Hehl, F. W., von der Heyde, P., Kerlick, G. D., On hypermomentum in general relativity III - Coupling hypermomentum to geometry, Zeitschrift für Naturforschung, Teil a, vol. 31a, July 1976, p. 823-827. DOI: 10.1515/zna-1976-0724
  • [22] F.W. Hehl, E.A. Lord, Yuval Ne’eman, Hadron Dilation, Shear and Spin as Components of the Intrinsic Hypermomentum Current and Metric Affine Theory of Gravitation, Phys.Lett.B 71 (1977) 432-434
  • [23] F. W. Hehl, J. Nitsch and P. von der Heyde, in General Relativity and Gravitation, ed. A. Held (Plenum Press, New York, 1980).
  • [24] F.W. Hehl, J.D. McCrea, E.W. Mielke, Y. Ne’eman, Metric-Affine Gauge Theory of Gravity: Field Equations, Noether Identities, World Spinors, and Breaking of Dilation Invariance, Phys. Rept. 258: 1-171, 1995. arXiv:gr-qc/9402012, https://doi.org/10.48550/arXiv.gr-qc/9402012
  • [25] Claire Rigouzzo and and Sebastian Zell, Coupling metric-affine gravity to the standard model and dark matter fermions, Phys. Rev. D 108, 124067, Dec 2023, DOI: https://doi.org/10.1103/PhysRevD.108.124067.
  • [26] Wheeler, James T. , Poincare gauge gravity from nonmetric gravity, arXiv:2407.13867 [gr-qc], https://doi.org/10.48550/arXiv.2407.13867
  • [27] Wheeler, James T., Sources for torsion in Poincaré gauge gravity, Eur. Phys. J. C 83, 665 (2023), https://doi.org/10.1140/epjc/s10052-023-11812-4, https://doi.org/10.48550/arXiv.2303.14921
  • [28] [28] D. S. Klemm and L. Ravera, Einstein manifolds with torsion and nonmetricity, Phys. Rev. D 101 (2020) no.4, 044011 doi:10.1103/PhysRevD.101.044011 [arXiv:1811.11458 [gr-qc]].
  • [29] David Hochberg, Gunter Plunien, Theory of matter in Weyl space-time, Phys.Rev.D 43 (1991) 3358-3367, DOI: 10.1103/PhysRevD.43.3358